# EEG-based Graph-guided Domain Adaptation for Robust Cross-Session Emotion Recognition

### v1

@@ Abstract
Accurate recognition of human emotional states is critical for effective human–machine interaction. Electroencephalography (EEG) offers a reliable source for emotion recognition due to its high temporal resolution and its direct reflection of neural activity. Nevertheless, variations across recording sessions present a major challenge for model generalization. To address this issue, we propose EGDA, a framework that reduces cross-session discrepancies by jointly aligning the global (marginal) and class-specific (conditional) distributions, while preserving the intrinsic structure of EEG data through graph regularization. Experimental results on the SEED-IV dataset demonstrate that EGDA achieves robust cross-session performance, obtaining accuracies of 81.22%, 80.15%, and 83.27% across three transfer tasks, and surpassing several baseline methods. Furthermore, the analysis highlights the Gamma frequency band as the most discriminative and identifies the central-parietal and prefrontal brain regions as critical for reliable emotion recognition.
@@ I Introduction
Among physiological signals, EEGhas emerged as one of the most informative sources for emotion recognition. EEGdirectly reflects neural activity with high temporal resolution and reflects the dynamics of cognitive and affective processes in the brain [4]. Despite promising progress, EEG-based emotion recognition faces critical challenges. A major difficulty arises from the non-stationarity of EEGsignals, as their distributions vary across recording sessions, subjects, and environmental conditions [5]. In particular, even for the same subject, data collected on different sessions may exhibit significant distributional shifts due to changes in mental state, electrode placement, or environmental noise [5],[6]. Consequently, models trained on one session often perform poorly when applied to another, limiting generalization. This challenge motivates the development of cross-session domain adaptation techniques that can transfer knowledge from a labeled source session to an unlabeled [7].
Several approaches focus on EEG-based emotion recognition using graph-based models to capture the structural relationships among samples [8]. However, these models do not employ transfer learning to address distribution mismatches across sessions or subjects. As a result, while they effectively model inter-sample similarity, they overlook aligning feature distributions between different domains.
To address these issues, we propose a framework termed EGDA. By integrating domain adaptation with graph-based regularization, EGDAsimultaneously reduces both marginal (feature-level) and conditional (class-conditional) distribution discrepancies while preserving the intrinsic geometric and semantic structures of EEGdata. Specifically, EGDAleverages an iterative pseudo-labeling strategy to refine class-level alignment and employs a Laplacian graph to maintain neighborhood consistency during the adaptation process. This unified framework enables robust cross-session emotion recognition performance.
@@
  • We formulate a graph-guided domain adaptation framework for cross-session EEG emotion recognition, in which structural relationships among EEG samples are explicitly incorporated into the adaptation objective to preserve neighborhood consistency across sessions.
  • We introduce a source-domain within-class scatter minimization constraint to preserve class discriminability and stabilize the adaptation process under cross-session distribution shifts.
  • We formulate a unified subspace learning model with graph-based regularization that jointly aligns marginal and conditional distributions; the conditional alignment is iteratively refined via pseudo-label updates regularized by neighborhood structure and within-class scatter constraints, without requiring labeled target data.
@@ II Related Work
This section reviews two categories of closely related studies: (1) EEG-based emotion recognition models evaluated on the SEED-IV dataset, and (2) transfer learning and domain adaptation approaches relevant to the proposed EGDAframework. Since EGDAis a domain adaptation model, both categories are relevant for positioning our contributions.
@@ II-A SEED-IV–Based emotion recognition methods
GFILproposes a unified importance learning framework for EEG-based emotion recognition by jointly performing emotion classification and feature importance estimation within a least-squares regression formulation [9]. By introducing an auto-weighting mechanism, GFILenables to automatically learn the relative contributions of different EEG feature dimensions, providing both improved recognition performance and enhanced interpretability.
@@
JSFAintroduces a joint sample and feature importance assessment framework for EEG-based emotion recognition by integrating self-paced learning with feature self-weighting into a least-squares regression model [10]. The self-paced mechanism adaptively down-weights low-quality or noisy EEG samples.
Graph-based semi-supervised learning has been widely explored to exploit the intrinsic structural relationships among EEG samples. OGSSLproposes a unified framework that jointly learns an optimal graph structure and infers emotional labels for unlabeled EEG samples, enabling mutual reinforcement between graph construction and semi-supervised classification [11].
@@
SRAGLproposes a semi-supervised regression framework with adaptive graph learning for cross-session EEG-based emotion recognition [12]. It jointly estimates the emotional labels of unlabeled samples and learns a sample similarity graph within a unified optimization process, allowing the graph structure and label inference to mutually reinforce each other. By iteratively updating the regression model and graph adjacency matrix, SRAGL aims to capture the intrinsic relationships among EEG samples and improve emotion recognition performance across sessions.
@@
GASDLproposes a graph-adaptive semi-supervised discriminative subspace learning framework for EEG-based emotion recognition [13]. It jointly learns a discriminative subspace by minimizing intra-class scatter while preserving local sample relationships through an adaptively constructed maximum-entropy graph. The graph learning, subspace projection, and pseudo-label estimation are tightly coupled within a unified optimization process, enabling mutual reinforcement between structural modeling and label inference. However, the iterative updating of neighborhood relationships may introduce graph instability, particularly in noisy or cross-session EEG scenarios.
CPTMLproposes a coupled projection transfer metric learning framework for cross-session EEG-based emotion recognition, aiming to address both cross-session and cross-trial distribution discrepancies [14]. It jointly integrates domain alignment and graph-based metric learning into a unified optimization objective, where domain alignment extracts domain-invariant features while metric learning enforces discriminative relationships among EEG samples belonging to the same emotional state. By coupling these two components, CPTMLenables mutual reinforcement between data alignment and similarity modeling, leading to improved recognition performance across sessions. Nevertheless, the coupled optimization procedure can be computationally demanding, which may affect scalability in practical EEG applications.
@@
JAGPaddresses cross-subject EEG emotion recognition by unifying domain-invariant feature adaptation, graph-adaptive label propagation, and emotion estimation within a single objective function [15]. The graph structure is iteratively updated based on the shared subspace representations, allowing label propagation and feature alignment to mutually reinforce each other. However, as both graph construction and feature adaptation depend on intermediate pseudo-labels, errors in label estimation may propagate through iterations and affect the robustness of the learned representations.
@@ II-B Transfer Learning Approaches
TCAis a feature-based unsupervised domain adaptation method that learns a shared subspace by minimizing the discrepancy between source and target marginal distributions using MMD[16]. By projecting data into transfer components, TCA reduces domain shift at the distribution level. However, TCAdoes not explicitly account for class-conditional distribution differences and does not exploit source label information, which may limit its effectiveness when conditional distributions vary significantly across domains.
JDAextends TCAby explicitly aligning both marginal and conditional distributions between source and target domains within a unified subspace learning framework [17]. It leverages pseudo-labels for the target domain to estimate class-conditional discrepancies and iteratively refines the shared representation. While JDA achieves improved adaptation performance, its reliance on pseudo-label quality may affect robustness under severe distribution shifts.
VDAfurther addresses large domain shifts by jointly reducing marginal and conditional distribution discrepancies while promoting class discrimination through domain-invariant clustering [18]. By integrating distribution adaptation with within-class scatter minimization, VDAlearns a discriminative and transferable representation.
Collectively, these methods demonstrate that distribution alignment effectively reduces inter-session variability. Building on this foundation, our work integrates robust graph learning with transfer alignment to better preserve discriminative features.
@@ III-A Preliminary
Let $\mathcal{D}_{s}=\{\boldsymbol{X}_{s},\boldsymbol{Y}_{s}\}$ denote the labeled source domain with $n_{s}$ samples, and $\mathcal{D}_{t}=\{\boldsymbol{X}_{t}\}$ denote the unlabeled target domain with $n_{t}$ samples. Here, $\boldsymbol{X}_{s}\in\mathbb{R}^{d\times n_{s}}$ and $\boldsymbol{X}_{t}\in\mathbb{R}^{d\times n_{t}}$ represent the EEG feature matrices of the source and target domains, respectively, where $d$ is the feature dimensionality. The label matrix of the source domain is denoted by $\boldsymbol{Y}_{s}\in\mathbb{R}^{C\times n_{s}}$, with $C$ being the number of emotion classes and each column of $\boldsymbol{Y}_{s}$ is a one-hot encoded label vector corresponding to a source sample.
We further define the combined feature matrix as $\boldsymbol{X}=[\boldsymbol{X}_{s},\boldsymbol{X}_{t}]\in\mathbb{R}^{d\times n},$ where $n=n_{s}+n_{t}$ is the total number of samples.
@@
We assume that the source and target domains share the same feature space and label space, i.e., $\mathcal{X}_{s}=\mathcal{X}_{t}$ and $\mathcal{Y}_{s}=\mathcal{Y}_{t}$. Here, $\mathcal{X}$ denotes the $d$-dimensional EEG feature space, and $\mathcal{Y}$ denotes the label space. However, due to domain shift, the source and target domains follow different data distributions. Specifically, their marginal distributions over the feature space and conditional distributions over the labels may differ, i.e., $P_{s}(x)\neq P_{t}(x)$ and $P_{s}(y\mid x)\neq P_{t}(y\mid x)$.
@@
To mitigate distribution discrepancy between domains, a linear projection matrix $\boldsymbol{A}\in\mathbb{R}^{d\times d_{f}}$ is learned, where $d$ and $d_{f}\leq d$ denote the original and projected feature dimensionalities, respectively. The projection $\boldsymbol{A}^{T}$ maps the source and target features to a shared subspace, yielding $\boldsymbol{A}^{T}\boldsymbol{X}_{s}$ and $\boldsymbol{A}^{T}\boldsymbol{X}_{t}$. This shared representation facilitates better alignment of marginal and conditional distributions while preserving discriminative information for emotion recognition.
III-B Problem Statement
Given a labeled source-domain EEG dataset $\mathcal{D}_{s}=\{\boldsymbol{X}_{s},\boldsymbol{Y}_{s}\}$ and an unlabeled target-domain EEG dataset $\mathcal{D}_{t}=\{\boldsymbol{X}_{t}\}$ collected from different sessions, the goal of this work is to learn a transferable model that can accurately recognize emotional states in the target domain. Due to distribution discrepancies across sessions, models trained on the source domain often generalize poorly to the target domain. Moreover, the absence of target labels makes direct supervised learning infeasible. Therefore, the key challenge is to jointly reduce cross-session distribution mismatch while inferring reliable target-domain labels, without access to labeled target data.
III-C Model Formulation
@@
In this paper, we propose a joint adaptation framework that learns a low-dimensional representation $\boldsymbol{A}\in\mathbb{R}^{d\times d_{f}}$ and simultaneously improving the target-domain labels. Since the target domain is unlabeled (i.e., $P_{t}(\boldsymbol{Y}_{t}|\boldsymbol{X}_{t})$ cannot be directly observed), we employ an EM-like strategy that iteratively updates both the the projection matrix and the target pseudo-labels..
@@
A central goal of transfer learning is to reduce distributional divergence between source and target domains. In EEG-based emotion recognition, session variability, noise, and experimental conditions cause distribution shifts, which degrade the performance of classifiers trained solely on source data. To address this, EGDAlearns a linear projection matrix $\boldsymbol{A}$ that maps both domains into a shared subspace where their marginal distributions are close. This principle can be implemented by the following objective function:
@@
Aligning only marginal distributions does not guarantee semantic consistency across domains, since it ignores class information and can lead to samples from different classes being mapped close to each other in the latent space. Consequently, class-conditional misalignment may still exist, especially in noisy EEGdata. To address this issue, EGDAaligns conditional distributions for each class using pseudo-labels for target samples.
The class-wise alignment is defined as:
$\min_{\boldsymbol{A}}\sum_{c=1}^{C}\left\|\frac{1}{n_{s}^{(c)}}\sum_{\boldsymbol{x}_{i}\in\boldsymbol{X}_{s}^{(c)}}\boldsymbol{A}^{T}\boldsymbol{x}_{i}-\frac{1}{n_{t}^{(c)}}\sum_{\boldsymbol{x}_{j}\in\boldsymbol{X}_{t}^{(c)}}\boldsymbol{A}^{T}\boldsymbol{x}_{j}\right\|^{2}$ (4)
@@
This can be expressed in matrix form as:
@@
$(\boldsymbol{M}_{c})_{ij}=\begin{cases}\frac{1}{(n_{l}^{{(c)}})^{2}}&\boldsymbol{x}_{i},\boldsymbol{x}_{j}\in\boldsymbol{X}_{l}^{{(c)}}\\ \frac{1}{(n_{t}^{(c)})^{2}}&\boldsymbol{x}_{i},\boldsymbol{x}_{j}\in\boldsymbol{X}_{t}^{(c)}\\ -\frac{1}{n_{l}^{(c)}n_{t}^{(c)}}&\boldsymbol{x}_{i}\in\boldsymbol{X}_{l}^{{(c)}},\ \boldsymbol{x}_{j}\in\boldsymbol{\boldsymbol{X}}_{t}^{(c)}\\ -\frac{1}{n_{l}^{{(c)}}n_{t}^{{(c)}}}&\boldsymbol{x}_{j}\in\boldsymbol{X}_{l}^{(c)},\ \boldsymbol{x}_{i}\in\boldsymbol{X}_{t}^{(c)}\\ 0&\text{otherwise}\end{cases}$ (6)
To enhance separability in the latent space, samples belonging to the same class in the source domain should form compact clusters. This property can be enforced by minimizing the within-class scatter matrix, defined as:
@@
$\boldsymbol{\mu}_{c}=\frac{1}{n_{s}^{(c)}}\sum_{\boldsymbol{x}_{i}\in\boldsymbol{X}_{s}^{(c)}}\boldsymbol{x}_{i},$ (8)
and $n_{s}^{(c)}$ denotes the number of source samples in class $c$.
To encourage the compactness of intra-class representations, EGDAformulates the following objective:
@@
$\min_{s_{ij}\geq 0,\;\boldsymbol{s}^{i}\boldsymbol{1}=1}\sum_{i,j=1}^{n}\|\boldsymbol{x}_{i}-\boldsymbol{x}_{j}\|_{2}^{2}s_{ij},$ (10)
@@
However, problem (10) may lead to a trivial solution in which only one element of $\boldsymbol{s}^{i}$ equals one while all others are zero [19]. Although this minimizes the objective, it produces a degenerate graph where each sample is connected only to its nearest neighbor. This fails to capture the local manifold structure and undermines graph-based regularization. To address this, a regularization term is introduced:
$\min_{\boldsymbol{S}}\sum_{i,j=1}^{n}\Big(\|\boldsymbol{x}_{i}-\boldsymbol{x}_{j}\|_{2}^{2}s_{ij}+\gamma s_{ij}^{2}\Big).$ (11)
@@
$\min_{\boldsymbol{S}}\sum_{j=1}^{n}\Big(\|\boldsymbol{x}_{i}-\boldsymbol{x}_{j}\|_{2}^{2}s_{ij}+\gamma s_{ij}^{2}\Big).$ (12)
@@
$\min_{\boldsymbol{S}}\Big\|\boldsymbol{s}^{i}+\frac{\boldsymbol{m}^{i}}{2\gamma}\Big\|_{2}^{2},$ (13)
@@
The first term of (14) can be written in matrix form as [9]:
@@
To prevent overfitting and ensure generalizable solutions, Frobenius norm regularization is applied:
@@
This penalizes large values in the projection matrix, encouraging smoother, more robust representations and avoiding fitting noise in the EEG data.
@@
$+\mu\,\operatorname{Tr}\!\big(\boldsymbol{A}^{T}\boldsymbol{X}\boldsymbol{L}_{s}\boldsymbol{X}^{T}\boldsymbol{A}\big)+\lambda\,\operatorname{Tr}\!\big(\boldsymbol{A}^{T}\boldsymbol{A}\big)+\gamma\,\operatorname{Tr}\!\big(\boldsymbol{S}^{T}\boldsymbol{S}\big)$
@@
where $\boldsymbol{H}=\boldsymbol{I}-\frac{1}{n}\boldsymbol{11}^{T}$ is the centering matrix, and $\alpha$, $\beta$, $\mu$, $\lambda$, and $\gamma$ are trade-off parameters controlling distribution alignment, within-class compactness, graph regularization, projection regularization, and similarity smoothness, respectively.
In the experiments, the values of the coefficients $\alpha$, $\beta$, $\gamma$, $\mu$, and $\lambda$ were selected randomly from the set $\{0.001,0.01,0.1,0.2,0.5,1,2,5,10,50,100$}\} to adjust the relative importance of the corresponding terms in the objective function.
III-D Model Optimization
The objective function of EGDAdepends on two variables, the projection matrix $\boldsymbol{A}$ and the similarity matrix $\boldsymbol{S}$. We optimize this objective using the alternating direction method, updating one variable while keeping the other fixed. The detailed update rules are as follows:
(1) $\boldsymbol{S}$ Optimziation:

### v2

@@ Abstract
Accurate recognition of human emotional states is critical for effective human–machine interaction. Electroencephalography (EEG) offers a reliable source for emotion recognition due to its high temporal resolution and direct reflection of neural activity. Nevertheless, variations across recording sessions present a major challenge for model generalization. To address this issue, we propose EEG-based graph-guided Domain adaptation (EGDA), a framework that reduces cross-session discrepancies by jointly aligning marginal and class-conditional distributions, while preserving the intrinsic structure of EEG data through graph regularization. Experimental results on the SEED-IV dataset demonstrate robust cross-session performance, achieving accuracies of 81.22%, 80.15%, and 83.27% across three transfer tasks and outperforming several baseline methods. Furthermore, the analysis highlights the Gamma frequency band as the most discriminative and identifies central-parietal and prefrontal brain regions as critical for reliable emotion recognition.
@@ I Introduction
Among physiological signals, electroencephalography (EEG) has emerged as one of the most informative sources for emotion recognition. EEG directly reflects neural activity with high temporal resolution and captures the dynamics of cognitive and affective processes in the brain [4]. Despite promising progress, EEG-based emotion recognition faces critical challenges. A major difficulty arises from the non-stationarity of EEG signals, as their distributions vary across recording sessions, subjects, and environmental conditions [5]. In particular, even for the same subject, data collected on different sessions may exhibit significant distributional shifts due to changes in mental state, electrode placement, or environmental noise [5, 6]. Consequently, models trained on one session often perform poorly when applied to another, limiting generalization. This challenge motivates the development of cross-session domain adaptation techniques that can transfer knowledge from a labeled source session to an unlabeled target session [7].
Several approaches focus on EEG-based emotion recognition using graph-based models to capture the structural relationships among samples [8]. However, these models typically do not employ transfer learning to address distribution mismatches across sessions or subjects. As a result, while they effectively model inter-sample similarity, they overlook aligning feature distributions between different domains.
To address these challenges, we propose a unified framework termed EGDA, which integrates domain adaptation and graph-based regularization for cross-session EEG-based emotion recognition. This proposed framework simultaneously mitigates discrepancies in the marginal distributions and class-conditional distributions while preserving the intrinsic geometric and semantic structures of EEG data. Specifically, EGDA employs an iterative pseudo-labeling strategy to progressively refine class-level alignment and incorporates a Laplacian graph regularizer to maintain neighborhood consistency throughout the adaptation process. As a result, the proposed framework enables robust and effective emotion recognition performance across different recording sessions.
@@
  • A graph-guided domain adaptation framework is proposed for cross-session EEG emotion recognition, in which graph regularization is explicitly integrated into the distribution alignment process to preserve local neighborhood consistency and to mitigate pseudo-label instability across sessions.
  • A source-domain within-class scatter minimization constraint is introduced to stabilize conditional distribution alignment by enforcing compact and discriminative class structures, thereby reducing error propagation during iterative pseudo-label refinement.
  • A unified subspace learning model is developed to jointly align marginal and class-conditional distributions, where conditional alignment is iteratively refined through pseudo-label updates under the joint regularization of neighborhood structure and within-class scatter, without requiring labeled target data.
@@ II Related Work
This section reviews two categories of closely related studies: (1) EEG-based emotion recognition models evaluated on the SEED-IV dataset, and (2) transfer learning and domain adaptation approaches relevant to the proposed EGDA framework. Since EGDA is a domain adaptation model, both categories are essential for positioning our contributions.
@@ II-A SEED-IV–Based Emotion Recognition Methods
Graph-regularized feature importance learning (GFIL) introduces a unified importance learning framework that jointly performs emotion classification and EEG feature importance estimation within a graph-regularized least-squares regression model [9]. By incorporating an auto-weighting variable, GFIL adaptively assigns different contributions to EEG feature dimensions, enabling simultaneous feature ranking, frequency-band importance analysis, and channel selection. The graph regularization term further enforces label consistency among samples sharing the same emotional state, resulting in improved recognition accuracy and enhanced interpretability in cross-session settings.
@@
Building upon the idea of adaptive weighting, Joint sample and feature importance assessment (JSFA) proposes a joint sample and feature importance assessment framework that integrates self-paced learning with feature self-weighting into a least-squares regression formulation [10]. By progressively emphasizing high-quality EEG samples and down-weighting noisy or ambiguous ones, JSFA improves robustness against data variability and enhances emotion recognition performance.
From a semi-supervised perspective, Optimal graph-coupled semi-supervised learning (OGSSL) presents an optimal graph-coupled semi-supervised learning framework that jointly learns an adaptive similarity graph and infers labels for unlabeled EEG samples within a unified optimization objective [11]. The mutual reinforcement between graph construction and label propagation allows OGSSL to effectively exploit both labeled and unlabeled data for emotion recognition.
@@
Similarly, Semi-supervised regression with adaptive graph learning (SRAGL) proposes a semi-supervised regression framework with adaptive graph learning for cross-session EEG-based emotion recognition [12]. By formulating emotion recognition as a semi-supervised regression problem and iteratively updating the sample similarity graph, SRAGL allows the regression model and graph structure to mutually reinforce each other, leading to improved generalization across sessions.
@@
Graph adaptive semi-supervised discriminative subspace learning (GASDSL) introduces a graph-adaptive semi-supervised discriminative subspace learning framework that learns a low-dimensional discriminative representation while preserving local sample relationships through an adaptively constructed maximum-entropy graph [13]. Although GASDSL effectively integrates graph learning, subspace projection, and pseudo-label estimation, the tight coupling among these components may lead to instability, particularly in noisy or cross-session EEG scenarios.
To explicitly address cross-session distribution discrepancies, Coupled projection transfer metric learning (CPTML) proposes a coupled projection transfer metric learning framework that jointly performs domain alignment and graph-based metric learning [14]. By enforcing discriminative relationships among EEG samples through intrinsic and penalty graphs while simultaneously aligning feature distributions, CPTML achieves improved cross-session emotion recognition performance. However, the coupled optimization strategy and metric learning component introduce considerable computational complexity, which may limit scalability in large-scale EEG applications.
@@
Joint adaptation with graph propagation (JAGP) focuses on cross-subject EEG emotion recognition by unifying domain-invariant feature adaptation, graph-adaptive label propagation, and emotion estimation within a single objective function [15]. The graph structure is iteratively updated based on shared subspace representations, enabling label propagation and feature adaptation to mutually reinforce each other. Nevertheless, since both graph construction and feature adaptation rely on intermediate pseudo-labels, error accumulation during iterations may adversely affect robustness.
Although the above methods have demonstrated strong performance on the SEED-IV dataset, their modeling strategies differ fundamentally from unsupervised domain adaptation frameworks designed for cross-session transfer. Compared with existing graph-based domain adaptation methods, EGDA exhibits several key distinctions. Unlike CPTML, which tightly couples domain alignment with explicit graph-based metric learning in both source and target domains, EGDA adopts a lightweight subspace learning formulation with Laplacian regularization and avoids metric learning, resulting in improved computational efficiency and stability. In contrast to JAGP, EGDA decouples graph modeling from target label estimation and stabilizes conditional distribution alignment by incorporating within-class scatter information exclusively from the source domain. Moreover, unlike semi-supervised graph-based approaches such as GASDSL, EGDA explicitly aligns both marginal and class-conditional distributions, making it more suitable for unsupervised cross-session EEG emotion recognition.
@@ II-B Transfer Learning Approaches
Transfer component analysis (TCA) is an unsupervised domain adaptation method that learns a shared subspace by minimizing the discrepancy between source and target marginal distributions using maximum mean discrepancy (MMD) [16]. By projecting data into transfer components, TCA reduces domain shift at the distribution level. However, TCA does not explicitly account for class-conditional distribution differences and does not exploit source label information, which may limit its effectiveness when conditional distributions vary significantly across domains.
Joint distribution adaptation (JDA) extends TCA by explicitly aligning both marginal and conditional distributions between source and target domains within a unified subspace learning framework [17]. It leverages pseudo-labels for the target domain to estimate class-conditional discrepancies and iteratively refines the shared representation.
Visual domain adaptation (VDA) further addresses large domain shifts by jointly reducing marginal and conditional distribution discrepancies while promoting class discrimination through domain-invariant clustering [18]. By integrating distribution adaptation with within-class scatter minimization, VDA learns a discriminative and transferable representation.
Collectively, these methods demonstrate that distribution alignment effectively reduces inter-session variability. Building on this foundation, our work integrates robust graph learning with transfer alignment to better preserve discriminative structures during cross-session adaptation.
@@ III-A Problem Statement
Let $\mathcal{D}_{s}=\{\boldsymbol{X}_{s},\boldsymbol{Y}_{s}\}$ denote a labeled source-domain EEG dataset with $n_{s}$ samples, and $\mathcal{D}_{t}=\{\boldsymbol{X}_{t}\}$ denote an unlabeled target-domain EEG dataset with $n_{t}$ samples. Here, $\boldsymbol{X}_{s}\in\mathbb{R}^{d\times n_{s}}$ and $\boldsymbol{X}_{t}\in\mathbb{R}^{d\times n_{t}}$ represent the EEG feature matrices of the source and target domains, respectively, where $d$ is the feature dimensionality. We further define the combined feature matrix as $\boldsymbol{X}=[\boldsymbol{X}_{s},\boldsymbol{X}_{t}]\in\mathbb{R}^{d\times n},$ where $n=n_{s}+n_{t}$ is the total number of samples. The source-domain label matrix is denoted by $\boldsymbol{Y}_{s}\in\mathbb{R}^{C\times n_{s}}$, where $C$ is the number of emotion classes (e.g., sadness, fear, happiness, and neutrality), and each column of $\boldsymbol{Y}_{s}$ is a one-hot encoded label vector corresponding to a source sample.
@@
We assume that the source and target domains share the same feature space and label space, i.e., $\mathcal{X}_{s}=\mathcal{X}_{t}$ and $\mathcal{Y}_{s}=\mathcal{Y}_{t}$, but follow different data distributions due to cross-session domain shift. In particular, both the marginal feature distributions and the conditional label distributions may differ, i.e., $P_{s}(x)\neq P_{t}(x)$ and $P_{s}(y\mid x)\neq P_{t}(y\mid x)$.
@@
To reduce the distribution discrepancy between domains, a linear projection matrix $\boldsymbol{A}\in\mathbb{R}^{d\times d_{f}}$ is learned to map both source and target EEG features into a shared subspace, yielding $\boldsymbol{A}^{T}\boldsymbol{X}_{s}$ and $\boldsymbol{A}^{T}\boldsymbol{X}_{t}$. This shared representation aims to align the source and target distributions while preserving discriminative information for emotion recognition.
Under this setting, we address a multi-class EEG-based emotion recognition problem without access to labeled target data. The objective is to learn a transferable model that can accurately recognize emotional states in the target domain despite cross-session distribution mismatch.
III-B Model Formulation
@@
In this paper, we propose a joint adaptation framework that learns a low-dimensional representation $\boldsymbol{A}\in\mathbb{R}^{d\times d_{f}}$ and simultaneously improves the target-domain labels. Since the target domain is unlabeled (i.e., $P_{t}(\boldsymbol{Y}_{t}|\boldsymbol{X}_{t})$ cannot be directly observed), we employ an expectation–maximization (EM)-like strategy that iteratively updates both the projection matrix and the target pseudo-labels.
@@
A central goal of transfer learning is to reduce distributional divergence between source and target domains. In EEG-based emotion recognition, session variability, noise, and experimental conditions cause distribution shifts, which degrade the performance of classifiers trained solely on source data. To address this, EGDA learns a linear projection matrix $\boldsymbol{A}$ that maps both domains into a shared subspace where their marginal distributions are close. This objective can be formulated as
@@
Aligning only marginal distributions does not guarantee semantic consistency across domains, since it ignores class information and can lead to samples from different classes being mapped close to each other in the latent space. Consequently, class-conditional misalignment may still exist. To address this issue, EGDA aligns conditional distributions for each class using pseudo-labels for target samples.
Specifically, the class-wise conditional alignment is formulated as
$\min_{\boldsymbol{A}}\sum_{c=1}^{C}\left\|\frac{1}{n_{s}^{(c)}}\sum_{\boldsymbol{x}_{i}\in\boldsymbol{X}_{s}^{(c)}}\boldsymbol{A}^{T}\boldsymbol{x}_{i}-\frac{1}{n_{t}^{(c)}}\sum_{\boldsymbol{x}_{j}\in\boldsymbol{X}_{t}^{(c)}}\boldsymbol{A}^{T}\boldsymbol{x}_{j}\right\|^{2}.$ (4)
@@
This objective can be equivalently expressed in matrix form as
@@
$(\boldsymbol{M}_{c})_{ij}=\begin{cases}\frac{1}{(n_{s}^{{(c)}})^{2}}&\boldsymbol{x}_{i},\boldsymbol{x}_{j}\in\boldsymbol{X}_{s}^{{(c)}}\\ \frac{1}{(n_{t}^{(c)})^{2}}&\boldsymbol{x}_{i},\boldsymbol{x}_{j}\in\boldsymbol{X}_{t}^{(c)}\\ -\frac{1}{n_{s}^{(c)}n_{t}^{(c)}}&\boldsymbol{x}_{i}\in\boldsymbol{X}_{s}^{{(c)}},\ \boldsymbol{x}_{j}\in\boldsymbol{X}_{t}^{(c)}\\ -\frac{1}{n_{s}^{{(c)}}n_{t}^{{(c)}}}&\boldsymbol{x}_{j}\in\boldsymbol{X}_{s}^{(c)},\ \boldsymbol{x}_{i}\in\boldsymbol{X}_{t}^{(c)}\\ 0&\text{otherwise}\end{cases}$ (6)
To enhance separability in the latent space, samples belonging to the same class in the source domain should form compact clusters. This property can be encouraged by minimizing the within-class scatter matrix, defined as:
@@
$\boldsymbol{\mu}_{c}=\frac{1}{n_{s}^{(c)}}\sum_{\boldsymbol{x}_{i}\in\boldsymbol{X}_{s}^{(c)}}\boldsymbol{x}_{i}.$ (8)
and $n_{s}^{(c)}$ represents the number of source-domain samples belonging to class $c$.
Based on this formulation, EGDA incorporates the following objective in the projected space to promote intra-class compactness:
@@
$\min_{\boldsymbol{S}}\sum_{i,j=1}^{n}\Big(\|\boldsymbol{x}_{i}-\boldsymbol{x}_{j}\|_{2}^{2}s_{ij}\Big),$ (10)
$\text{s.t.}\quad s_{ij}\geq 0,\;\boldsymbol{s}^{i}\boldsymbol{1}=1.$
@@
However, problem (10) may lead to a trivial solution in which only one element of $\boldsymbol{s}^{i}$ equals one while all others are zero [19]. Although this minimizes the objective, it produces a degenerate graph where each sample is connected only to its nearest neighbor. This fails to capture the local manifold structure of the data and consequently undermines the effectiveness of graph-based regularization.
To address this, a regularization term is introduced:
$\min_{\boldsymbol{S}}\sum_{i,j=1}^{n}\Big(\|\boldsymbol{x}_{i}-\boldsymbol{x}_{j}\|_{2}^{2}s_{ij}+\gamma s_{ij}^{2}\Big),$ (11)
@@
$\min_{\boldsymbol{S}}\sum_{j=1}^{n}\Big(\|\boldsymbol{x}_{i}-\boldsymbol{x}_{j}\|_{2}^{2}s_{ij}+\gamma s_{ij}^{2}\Big),$ (12)
@@
$\min_{\boldsymbol{s}^{i}}\Big\|\boldsymbol{s}^{i}+\frac{\boldsymbol{m}^{i}}{2\gamma}\Big\|_{2}^{2},$ (13)
@@
The first term of (14) can be written in matrix form as [9]
@@
To prevent overfitting and improve generalization, Frobenius norm regularization is applied to the projection matrix:
@@
This penalizes large values in the projection matrix, encouraging smoother and more robust representations while reducing sensitivity to noise in EEG data.
@@
$+\mu\,\operatorname{Tr}\!\big(\boldsymbol{A}^{T}\boldsymbol{X}\boldsymbol{L}_{s}\boldsymbol{X}^{T}\boldsymbol{A}\big)+\lambda\,\operatorname{Tr}\!\big(\boldsymbol{A}^{T}\boldsymbol{A}\big)+\gamma\,\operatorname{Tr}\!\big(\boldsymbol{S}^{T}\boldsymbol{S}\big),$
@@
where $\boldsymbol{H}=\boldsymbol{I}-\frac{1}{n}\boldsymbol{11}^{T}$ is the centering matrix. Here, $c=0$ corresponds to marginal distribution alignment, while $c=1,\dots,C$ denote class-conditional alignment terms. The parameters $\alpha$, $\beta$, $\mu$, $\lambda$, and $\gamma$ are trade-off coefficients controlling distribution alignment, within-class compactness, graph regularization, projection regularization, and similarity smoothness, respectively.
Due to the large search space, the hyperparameters were selected by randomly sampling from the predefined candidate set $\{10^{-3},10^{-2},10^{-1},2\times 10^{-1},5\times 10^{-1},10^{0},2\times 10^{0},5\times 10^{0},10^{1},5\times 10^{1},10^{2}\}$.
III-C Model Optimization
The objective function of EGDA depends on two variables, the projection matrix $\boldsymbol{A}$ and the similarity matrix $\boldsymbol{S}$. We optimize this objective using an alternating optimization strategy, updating one variable while keeping the other fixed.
(1) $\boldsymbol{S}$ Optimization: